Shift report 20010922 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Sep 22 2001 - 16:52:34 EDT

  • Next message: Jens Jørgen Gaardhøje: "Re: Run plan considerations."

    ------------------------------------------------------------
    Date: Sat Sep 22 16:52:34 2001, Shift: 08:00-16:00
    Supervisor: JN, with: FV
    Log book#9 pages 25 - 25
    
    Shift Summary:
    ------------------------------------------------------------
    Machine still recovering following the dip last night.
    No data runs. Just waiting.
    
    Add H1,H2 scalers #13,#14 on pet page
    H1 (slat 1-16 up/down coinc) sum
    H2  (slat 2-8 up/down coinc) sum
    
    Magents were restored by watch ~10.00 am.
    
    Notice the continuing lack of 4032 control -
    we should request a scehduled maint. this week for this.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Sep 22 2001 - 18:12:58 EDT